OpenStudy (anonymous):

Solve the following equation. log3(x-2)=-1

OpenStudy (anonymous):

rewrite in equivelant exponential from as \[3^{-1}=x-2\] and solve for x

OpenStudy (anonymous):

so would i subtract the -2 from 3

OpenStudy (anonymous):

it is not 3, it is \[3^{-1}=\frac{1}{3}\]

OpenStudy (anonymous):

\[\frac{1}{3}+2=\frac{7}{3}=x\]